A sample of nitrogen gas at a pressure of 0.744 atm and a temperature of 25.2 °C, occupies a volume of 11.6 liters. If the gas is allowed to expand at constant temperature to a volume of 16.1 liters, the pressure of the gas sample will be
_________ atm.
1
Expert's answer
2019-03-14T04:01:02-0400

P1V1 = P2V2

P2 = P1V1/V2

P2 = 0.744*11.6/16.1 = 0.536 (atm)